# DOEPatch: Dynamically Optimized Ensemble Model for Adversarial Patches Generation

### v1

@@ IV-B Digital Attack Evaluations
All adversarial patches were applied to the images in the Inria test set, and the effectiveness of the attacks was evaluated by calculating the changes in the AP of the person category in the output of different object detection models. Experimental results in the digital world are shown in Tab.I, where “Clean” represents the AP value of the person category in the original test set. The AP values for the test sets with 12 additional adversarial patches were tested using 5 mainstream object detection models. Under the detection of YOLOv2, DOEPatch(YY) reduced the AP value by 67.49%, slightly lower than AdvPatch. However, the attack performance of DOEPatch(YY) is better than other adversarial patches such as TCEGA. Meanwhile, the best attack performance was achieved under the detection of YOLOv3, where DOEPatch(YY) was able to reduce the AP value to 29.20%. Our energy-based adversarial patch EmPatch(Y2) achieved an AP reduction of 54.11% under the detection of YOLOv3, and its attack performance was far superior to the test results of other patches. For the more complex Faster-VGG detection algorithm, DOEPatch(YF) and DOEPatch(YYF) achieved the highest attack effect, reducing AP values by 43.94% and 44.81%, respectively. They also demonstrated strong attack effectiveness under the detection of YOLOv3, reducing AP values by 41.91% and 41.65%, respectively. Due to differences in the dataset, UpcPatch only reduced the AP value by 6.10% in the case of Faster R-CNN.
@@
To comprehensively compare the transferability of adversarial patches, we conducted black-box attack tests on YOLOv4 and Faster-Res detection algorithms. Our trained DOEPatch(YY) can reduce the AP value by 50.55% in the detection of YOLOv4, while AdvPatch and TCEGA can only reduce it by 29.14% and 28.21%, respectively. Under Faster-Res, the attack performance was generally poor, but DOEPatch(YY) and DOEPatch(YF) still achieved the best attack results, reducing the AP value by 18.55% and 17.05%, respectively. Overall, the adversarial patches generated based on a dynamically optimized ensemble model exhibited strong attack performance across multiple targeted models and demonstrated favorable transferability by causing significant output interference against unknown black-box models.
@@ IV-C Physical Attack Evaluations
To validate the attack efficacy of the adversarial patches in the physical world, we printed them on T-shirts to simulate the target detection process in real-world settings. Specifically, the adversarial patches in the digital world were composed of 300$\times$300 pixels, which were enlarged to 30cm×$\times$30cm and printed onto white cotton T-shirts using offset printing techniques for physical-world testing, as illustrated in Fig.6.
@@
To simulate realistic detection scenarios, we recorded two videos of persons’ movements using a c920e camera in indoor and outdoor environments, respectively. One of the persons wore a T-shirt with the adversarial patch, and the attack efficacy of the patch in the physical world is demonstrated in Fig.7. Both YOLOv2 and YOLOv3 were able to detect all persons in all frames without adversarial T-shirts, across different environments. However, using adversarial T-shirts created with DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF), another person could not be detected in both indoor and outdoor environments. According to the results, the adversarial patches generated were found to be effective in manipulating the recognition capability of object detection models in the physical world. Specifically, the patches were able to enable individuals wearing adversarial T-shirts to evade detection by object detection models, while also exhibiting robust transferability across multiple object detection models. These findings highlight the potential vulnerability of real-world object detection models to adversarial attacks and emphasize the need for further research to enhance the robustness of such models against potential threats.
@@
Figure 7: In the detection results of indoor and outdoor scenes, the adversarial patches printed on the front and back can affect the object detector recognition.
IV-D Interpretability of the Adversarial Patch
@@
Although most works have studied the interpretability of adversarial examples[50], there is no comprehensive explanation of the attack performance for adversarial patches. Consequently, we offer a more comprehensive and insightful explanatory analysis of the phenomenon of how adversarial patches perturb target detection predictions, utilizing the energy function perspective and the visual interpretation tool of Grad-CAM[30]. Adversarial patches generated by the energy function and the dynamically optimized ensemble model exhibit complex and distinct texture features in both color and feature distribution, resembling traditional adversarial patches. Fig.8(a) shows the detection result of DOEPatch(YYF) in YOLOv2, misidentified as the “teddy bear” category. Fig.8(b) depicts the detection result of DOEPatch(YF) in YOLOv3, misidentified as the “apple” category. Fig.8(c) shows the detection result of EmPatch(F) in Faster R-CNN, misidentified as the “potted plant” category. Merely analyzing partial detection results is insufficient for uncovering the underlying mechanisms of adversarial patches. From the perspective of energy function analysis, the training process reduces the energy to impair the object detection capability of the target detector on the “person” category by optimizing the adversarial patch. However, it is challenging to visualize the energy change directly, and we describe the energy change process by comparing the number of selected category candidate boxes after the screening. Fig.8(d) illustrates the changes in the candidate boxes’ output of YOLOv2 upon adding DOEPatch(YYF) to the test samples. Fig.8(e) shows the transformation of the candidate boxes’ output of YOLOv3 after adding DOEPatch(YF) to the test samples, and Fig.8(f) depicts the changes in the candidate boxes’ output of Faster R-CNN after adding EmPatch(F). The blue bars represent the number of candidate boxes output by the detection network after predicting the original samples, while the orange bars represent the output after detecting the adversarial samples. Specifically, the number of candidate boxes representing the “person” category sharply decreases after adding the adversarial patch, while some other categories’ quantities increase to a certain extent. By training adversarial patches with complicated features that enhance the energy of other categories, the energy of the person category is lowered to a certain extent.
Figure 8: (a)-(c) represent the misrecognition results of adversarial patches in object detection. (d)-(f) depict the results of changes in predicted categories.
@@
Merely analyzing from the energy perspective is not comprehensive enough. In order to further analyze the effects of the adversarial attack, we used the pre-trained VGG16 model on ImageNet and combined it with the Grad-CAM algorithm to demonstrate the changes in the attention maps of the target images before and after the attack. Two main changes in the results were observed from the experiments conducted on the test dataset. As shown in Fig.9, the (i) row represents the original sample and the adversarial sample after adding the adversarial patch, the (ii) row represents the recognition results under the detector, and the (iii) row represents the detection results under Grad-CAM. As shown in (ii), in Result 1, the person was not recognized as it was not attended to due to the attention being distracted by the adversarial patch. The feature distribution of the person was altered due to the presence of the adversarial patch. In Result 2, the adversarial patch was misrecognized, and all the attention of the model was attracted to the patch, leading to a change in the model’s decision. Overall, the adversarial patch has the ability to affect the distribution of the target features in the detection network, thereby interfering with the prediction results.
@@
IV-E Ablation Studies

### v2

@@ IV-B Digital Attack Evaluations
All adversarial patches were applied to the images in the Inria test set, and the effectiveness of the attacks was evaluated by calculating the changes in the AP of the person category in the output of different object detection models. Experimental results in the digital world are shown in Tab.I, where “Clean” represents the AP value of the person category in the original test set. The AP values for the test sets with 12 additional adversarial patches were tested using 5 mainstream object detection models. Under the detection of YOLOv2, DOEPatch(YY) reduced the AP value by 67.49%, slightly lower than AdvPatch. However, the attack performance of DOEPatch(YY) is better than other adversarial patches such as TCEGA. Meanwhile, the best attack performance was achieved under the detection of YOLOv3, where DOEPatch(YY) was able to reduce the AP value to 29.20%. Our energy-based adversarial patch EmPatch(Y2) achieved an AP reduction of 54.11% under the detection of YOLOv3, and its attack performance was far superior to the test results of other patches. For the more complex Faster-VGG detection algorithm, DOEPatch(YF) and DOEPatch(YYF) achieved the highest attack effects, reducing AP values by 43.94% and 44.81%, respectively. They also demonstrated strong attack effectiveness under the detection of YOLOv3, reducing AP values by 41.91% and 41.65%, respectively. Due to differences in the dataset, UpcPatch only reduced the AP value by 6.10% in the case of Faster R-CNN.
To comprehensively compare the transferability of adversarial patches, we conducted black-box attack tests on YOLOv4 and Faster-Res detection algorithms. Our trained DOEPatch(YY) can reduce the AP value by 50.55% in the detection of YOLOv4, while AdvPatch and TCEGA can only reduce it by 29.14% and 28.21%, respectively. Under Faster-Res, the attack performance was generally poor, but DOEPatch(YY) and DOEPatch(YF) still achieved the best attack results, reducing the AP value by 18.55% and 17.05%, respectively. The ensemble model we designed incorporates commonly used networks in object detection, such as ResNet, VGG, and others. During the training process, we ensured that adversarial patches could disrupt common feature extraction networks, thus affecting the recognition of object detectors, achieving a certain degree of generality. Moreover, the networks used in common target detectors will not be completely different. Therefore, it will also be effective in attacking against models other than the available ones. Overall, the adversarial patches generated based on a dynamically optimized ensemble model exhibited strong attack performance across multiple targeted models and demonstrated favorable transferability by causing significant output interference against unknown black-box models.
@@ IV-C Physical Attack Evaluations
To validate the attack efficacy of the adversarial patches in the physical world, we printed them on T-shirts to simulate the target detection process in real-world settings. Specifically, the adversarial patch in the digital world consists of 300$\times$300 pixels, which we enlarged to 30cm×$\times$30cm and printed the adversarial patch on a white cotton T-shirt via DTF-400A printer for testing in the physical world, as shown in Fig.6. Due to time and financial constraints, we have only produced three sets of counter patches, DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF). In order to reduce the pixel difference between the digital and physical worlds, based on the work of Mahmood Sharif et al.[38], we start by making a colour palette containing a one-fifth of the RGB colour. Then, after making it through this printer, the image is captured with the camera of the Redmi K50 Pro to obtain a set of RGB triplets printed in the physical world. In order to reduce the computational complexity, we convert this set of colour sets and into a set of 30 RGB triples that have the smallest change in distance from the original set of colours, i.e., we obtain the printable RGB triple C in Eq.7.
@@
In order to evaluate the difference in pixel values of the adversarial patch after printing, we evaluate it by calculating the histogram distribution of the hue channels in the HSV colour space. As shown in Fig.7, the image on the left is the adversarial patch after printing, and the curves on the right represent the variation of the patch’s hue distribution in the digital and physical worlds. We can find that the change curves in the digital and physical worlds basically overlap, and the similarity of the hue distributions in the adversarial patches DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF) are computed to be 0.8235, 0.8332, and 0.8207, respectively. In computing the histogram similarity, the range of the results is usually in [-1 ,1], with 1 indicating a perfect match and -1 indicating a complete mismatch. Therefore, we have greatly reduced the colour loss of digital to physical world transfer by the Non-Printable Score optimized adversarial patches.
Figure 7: The distribution of adversarial patch tones in both digital and physical worlds.
Regarding the smoothness of the adversarial patches, we directly calculate the loss of smoothness of the adversarial patches in the digital and physical worlds by using the Eq.8. The experimental results are shown in Tab.II. The overall smoothness of the adversarial patch in the physical world is lower than that in the digital world. The printer, due to the limitation of RGB colour space, some of the colours in the digital space cannot be printed out accurately, which makes the smoothness loss lower. The biggest decrease in the DOEPatch(YYF) reaches 0.00077, which also shows that the pixel values of the counterpatch after printing have very little change and have better smoothness.
TABLE II: Smoothness values of adversarial patches in the digital and physical worlds.
Digital Physical
Methods DOEPatch(YF) DOEPatch(YY) DOEPatch(YYF) DOEPatch(YF) DOEPatch(YY) DOEPatch(YYF)
$L_{smooth}$ 0.01333 0.01327 0.01332 0.01278 0.01307 0.01255
Due to time and venue constraints, we conducted physical experiments in four different scenarios, including indoor and outdoor environments, under various lighting conditions. In each scenario, two pedestrians of different body sizes were arranged. During each data collection session, one of the individuals wore the adversarial T-shirt. Firstly, to verify the transferability of adversarial patches, we collected four sets of videos for each adversarial patch within a distance range of 2 to 4 meters and an angle range of -45° to 45°. By extracting images from the videos, each adversarial patch obtained 160 test data samples. The experimental results, as shown in Tab.III, demonstrate that DOEPatch (YY) achieved the highest attack effectiveness in YOLOv2, reaching 56.87%. Our adversarial patches also demonstrated some attack effectiveness when confronted with the more complex Faster-VGG detection model.
TABLE III: ASR test results of adversarial patches under different detectors.
YOLOv2 YOLOv3 Faster-VGG YOLOv4 Faster-Res
DOEPatch(YY) 56.87 33.75 21.25 30.62 15.62
DOEPatch(YF) 35.00 28.12 13.75 25.00 8.75
DOEPatch(YYF) 43.75 20.62 16.87 21.87 6.87
@@
Partial prediction results as shown in Fig.8 indicate that in different environments, both YOLOv2 and YOLOv3 can detect all individuals not wearing adversarial T-shirts in images. However, they cannot detect another person wearing adversarial T-shirts crafted using DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF). According to the results, the adversarial patches generated were found to be effective in manipulating the recognition capability of object detection models in the physical world. Specifically, the patches were able to enable individuals wearing adversarial T-shirts to evade detection by object detection models, while also exhibiting robust transferability across multiple object detection models. These findings highlight the potential vulnerability of real-world object detection models to adversarial attacks and emphasize the need for further research to enhance the robustness of such models against potential threats.
@@
Figure 8: In the detection results in the four scenarios, the adversarial patches printed on the front and back can affect the object detector recognition.
IV-D Robustness Evaluations of the Adversarial Patch
To comprehensively assess the robustness of adversarial patches and their limitations in practical applications, we conducted tests based on two factors: distance and angle. Firstly, in each environment, we collected three datasets for each adversarial patch: short distance (2 to 4 meters), medium distance (5 to 7 meters), and long distance (8 to 10 meters), with 25 images per dataset. Therefore, each adversarial patch was validated on 100 images within each distance range. The experimental results, as shown in Fig.9, were verified using YOLOv2. DOEPatch(YY) achieved ASR of 53% at short distances, while DOEPatch(YYF) reached 44%, indicating good attack effectiveness. However, with increasing distance, the attack effectiveness of adversarial patches decreased at medium distances. At long distances, we observed that all adversarial patches failed to impact the target detector’s recognition. As the distance increased, the presentation of adversarial patches in the images became smaller and more blurred, leading to a decline in attack performance.
Figure 9: ASR test results of adversarial patches under different distances.
For the angle transformation experiment, we collected datasets within the range of 0° to 360° at a distance of two meters from the target for each adversarial patch. The ASR test results on YOLOv2 are shown in Fig.10. Each point of the marker was obtained by testing on 30 images. The adversarial patch DOEPatch(YY) had the best attack effect at 0° with an ASR of 70%. As the angle changes, the attack effectiveness of all adversarial patches declines. However, after 90°, the ASR values begin to rise again, reaching a peak at 180°. Observing the trend of the curves in the graph, we find that adversarial patches exhibit good attack effectiveness only when directly facing the camera. Furthermore, frontal testing yields significantly better results than testing from the rear. In side-facing scenarios, due to the limitations of the viewing angle, most of the area covered by the adversarial patches is not captured, leading to a substantial decrease in attack effectiveness. Distance and viewing angle interference with adversarial patches are unavoidable, and what we need to do is to minimize the impact of these factors as much as possible. Increasing the area of the adversarial patches and addressing blurriness issues can help mitigate the effects of distance. Regarding the influence of viewing angles, considering a full coverage approach can reduce the impact of occlusion to some extent. These issues mentioned above are what we need to address in our next steps.
Figure 10: ASR test results of adversarial patches under different angles.
To evaluate the effectiveness of our method on a new model, we retrained the adversarial patch DOEPatch against the more advanced YOLOv5s.pt open-source model. The YOLOv5[48] and YOLOv7[49] series of open-source models are more widely used in practical applications, with most real systems being based on these open-source models retrained to perform tasks. Therefore, we validated the attack effect of the adversarial patch on YOLOv5 and YOLOv7 models using the Inria test data, with the mAP testing results shown in Tab.IV. DOEPatch achieved a decrease of 39.2% in mAP on the YOLOv5s.pt model. Even on the complex YOLOv7x.pt model, the mAP was reduced from 95.2% to 59.4%, indicating that our attack method exhibits good robustness across different models.
TABLE IV: The mAP testing results of adversarial patches in different models.
YOLOv5s.pt YOLOv5n.pt YOLOv5x.pt YOLOv7.pt YOLOv7x.pt
Clean 0.934 0.921 0.959 0.957 0.952
DOEPatch 0.542 0.627 0.613 0.651 0.594
IV-E Interpretability of the Adversarial Patch
@@
Although most works have studied the interpretability of adversarial examples[50], there is no comprehensive explanation of the attack performance for adversarial patches. Consequently, we offer a more comprehensive and insightful explanatory analysis of the phenomenon of how adversarial patches perturb target detection predictions, utilizing the energy function perspective and the visual interpretation tool of Grad-CAM[30]. Adversarial patches generated by the energy function and the dynamically optimized ensemble model exhibit complex and distinct texture features in both color and feature distribution, resembling traditional adversarial patches. Fig.11(a1) shows the detection result of DOEPatch(YYF) in YOLOv2, misidentified as the “teddy bear” category. Fig.11(b1) depicts the detection result of DOEPatch(YF) in YOLOv3, misidentified as the “apple” category. Fig.11(c1) shows the detection result of EmPatch(F) in Faster R-CNN, misidentified as the “potted plant” category. Merely analyzing partial detection results is insufficient for uncovering the underlying mechanisms of adversarial patches. From the perspective of energy function analysis, the training process reduces the energy to impair the object detection capability of the target detector on the “person” category by optimizing the adversarial patch. However, it is challenging to visualize the energy change directly, and we describe the energy change process by comparing the number of selected category candidate boxes after the screening. As shown in the bar chart in Fig.11, the horizontal coordinates represent the predicted categories and the vertical coordinates represent the absolute number of candidate boxes of different categories output by the detector. Fig.11(a2) illustrates the changes in the candidate boxes’ output of YOLOv2 upon adding DOEPatch(YYF) to the test samples. Fig.11(b2) shows the transformation of the candidate boxes’ output of YOLOv3 after adding DOEPatch(YF) to the test samples, and Fig.11(c2) depicts the changes in the candidate boxes’ output of Faster R-CNN after adding EmPatch(F). The blue bars represent the number of candidate boxes output by the detection network after predicting the original samples, while the orange bars represent the output after detecting the adversarial samples. Specifically, the number of candidate boxes representing the “person” category sharply decreases after adding the adversarial patch, while some other categories’ quantities increase to a certain extent. By training adversarial patches with complicated features that enhance the energy of other categories, the energy of the person category is lowered to a certain extent.
Figure 11: (a1), (b1), (c1) denote the results of misidentification of adversarial patches in object detection. (a2), (b2), (c2) denote the results of the change in the absolute number of predicted category candidate boxes.
@@
Merely analyzing from the energy perspective is not comprehensive enough. In order to further analyze the effects of the adversarial attack, we used the pre-trained VGG16 model on ImageNet and combined it with the Grad-CAM algorithm to demonstrate the changes in the attention maps of the target images before and after the attack. Two main changes in the results were observed from the experiments conducted on the test dataset. As shown in Fig.12, the (i) row represents the original sample and the adversarial sample after adding the adversarial patch, the (ii) row represents the recognition results under the detector, and the (iii) row represents the detection results under Grad-CAM. As shown in (ii), in Result 1, the person was not recognized as it was not attended to due to the attention being distracted by the adversarial patch. The feature distribution of the person was altered due to the presence of the adversarial patch. In Result 2, the adversarial patch was misrecognized, and all the attention of the model was attracted to the patch, leading to a change in the model’s decision. The adversarial patch has the ability to affect the distribution of the target features in the detection network, thereby interfering with the prediction results.
@@
To provide a more comprehensive analysis of the effectiveness of adversarial patches, we propose a quantitative metric called the Disruption Area Ratio $R_{d}$. This metric aims to measure the proportion of different areas within the bounding box region of the target object, as determined by the true label, in the Grad-CAM outputs of both clean images and adversarial samples. Specifically, we first generate Grad-CAM heatmaps for both clean images and adversarial samples. Then, based on the true label, we identify the bounding box region of the target object in the heatmaps. Subsequently, we compute the proportion of different regions between the two heatmaps, which serves to evaluate the impact of adversarial samples on the bounding box region of the target object. We validated this metric on the Inria Person test set, computing the average across 288 test images. After adding the adversarial patch DOEPatch(YY), 60.92% of the regions in the target object heat map distribution are different from the original image. With the addition of DOEPatch(YF) and DOEPatch(YYF), results of 68.27% and 64.96% were also obtained, respectively. This indicates that the feature distribution of the target object has been severely disturbed, resulting in an incorrect recognition by the target detector.
IV-F Adversarial training Evaluations
Adversarial training entails introducing adversarial samples into the model, allowing it to learn features inherent in adversarial samples and thus bolstering the model’s resilience against adversarial attacks. To validate the effectiveness of our generated adversarial samples in fortifying the reliability of object detection models, we devised adversarial training experiments for evaluation. Firstly, we randomly extracted 120 images from the Inria Person dataset for the training set, 30 images for the validation set, and 180 images for the testing set. Then, we added adversarial patches DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF) to the extracted images, resulting in adversarial sample training, validation, and testing sets. We added the generated adversarial sample training set to the original training sets of YOLOv3 and Faster R-CNN, and retrained the models. Therefore, the adversarial training mode we designed is as follows:
  • YOLO-YY: Add adversarial samples generated by DOEPatch(YY) to the original data, then retrain YOLOv3.
  • YOLO-YF: Add adversarial samples generated by DOEPatch(YF) to the original data, then retrain YOLOv3.
  • YOLO-YYF: Add adversarial samples generated by DOEPatch(YYF) to the original data, then retrain YOLOv3.
  • FR-YY: Add adversarial samples generated by DOEPatch(YY) to the original data, then retrain Faster R-CNN.
  • FR-YF: Add adversarial samples generated by DOEPatch(YF) to the original data, then retrain Faster R-CNN.
  • FR-YYF: Add adversarial samples generated by DOEPatch(YYF) to the original data, then retrain Faster R-CNN.
We saved the model every 10 epochs and tested it on the validation set. After training for 100 epochs, we selected the model with the best results for subsequent testing. Tab.V shows the AP test results for all adversarial patches, where $\Delta$ represents the change in detection models after adversarial training compared to the original models. We observed significant improvements in the models after adversarial training, particularly in the testing set with adversarial samples. The YOLO-YY model achieved a 64.67% increase in AP on DOEPatch(YY), while the FR-YYF model demonstrated a 33.53% improvement on DOEPatch(YYF). Then, after adversarial training, the detection models exhibit strong resistance to interference from test data that possess features similar to the input adversarial samples. The YOLO-YF model achieved improvements of 48.01% and 40.21% on DOEPatch(YF) and DOEPatch(YYF) respectively. Meanwhile, the FR-YF model demonstrated enhancements of 33.93% and 34.09% on DOEPatch(YF) and DOEPatch(YYF) respectively. Furthermore, adversarial training not only enhances the model’s resistance to interference but also improves its performance on clean samples compared to the original model. YOLO-YF exhibited a 3.93% improvement on clean samples compared to the original model. Additionally, adversarial training does not necessarily grant the model stronger resistance to all adversarial patches. In the case of UpcPatch, the AP decreased in all tests involving YOLOv3 models. UpcPatch already yields high AP results when tested on the original model, making further enhancements challenging. Moreover, it possesses significant differences from the features of input adversarial samples, causing the model to lean towards recognizing the features of the input adversarial samples during training. In conclusion, although adversarial training may not render the model completely immune to adversarial patches, it can greatly enhance the robustness of detection models.
According to the experimental results of adversarial training, while training the detection model using adversarial samples as input data effectively improves performance, it still exhibits instability when faced with other unknown attacks. Some of the existing work on improving the adversarial training strategy based on adversarial samples to minimize the distance between clean and perturbed images, such as stylized adversarial training(SAT)[51], attribute-guided adversarial training[52]. Furthermore, the primary reason adversarial patches impact object detectors is their intricate surface texture and color distribution, which significantly influences the detector’s ability to extract image features. Therefore, some preprocessing operations can be added to the input data to reduce the effect of the antagonistic samples, such as adding a denoising module[53, 54] to filter the antagonistic noise. On the other hand, the vulnerability of the model itself is also an issue that cannot be ignored. The distillation-based[55, 56] approach compresses and transfers the knowledge from a complex large-scale model or an integrated model consisting of several models to a small model, which largely improves the stability of the small model. Hence, adversarial attack methods provide a new perspective for assessing the efficacy of models, emphasizing the significance of a model’s robustness against unknown variations. Our approach enables a deeper investigation into the vulnerabilities of object detection models, offering novel directions for enhancing the security of object detection models.
TABLE V: The AP test results of detection models after adversarial training on adversarial patches are presented. The left axis lists all adversarial patches, while the top axis lists the detection models. $\Delta$ indicates the change in AP compared to the original models after adversarial training.
YOLOv3 YOLO-YY($\Delta$) YOLO-YF($\Delta$) YOLO-YYF($\Delta$) FR FR-YY($\Delta$) FR-YF($\Delta$) FR-YYF($\Delta$)
Clean 89.25 92.18(2.93) 93.18(3.93) 90.41(1.16) 92.77 92.85(0.08) 93.32(0.55) 93.36(0.59)
AdvPatch 71.21 87.71(16.5) 88.68(17.47) 88.25(17.04) 70.3 85.34(15.04) 82.02(11.72) 83.26(12.96)
TCEGA 55.63 66.25(10.62) 88.37(32.74) 89.34(33.71) 74.65 83.65(9.00) 90.36(15.71) 89.11(14.46)
NaPatch 76.89 76.14(-0.75) 78.48(1.59) 78.97(2.08) 88.75 90.54(1.79) 90.46(1.71) 90.63(1.88)
EmPatch(Y1) 61.45 89.98(28.53) 75.13(13.68) 77.46(16.01) 64.38 87.77(23.39) 72.74(8.36) 79.67(15.29)
EmPatch(Y2) 30.87 62.75(31.88) 83.24(52.37) 81.24(50.37) 90.37 90.30(-0.07) 93.66(3.29) 90.77(0.40)
DOEPatch(YY) 26.17 87.84(61.67) 59.47(33.3) 63.35(37.18) 64.82 87.50(22.68) 75.14(10.32) 80.77(15.95)
EmPatch(F) 83.77 81.98(-1.79) 86.91(3.14) 87.91(4.14) 48.92 58.47(9.55) 80.82(31.90) 77.96(29.04)
UpcPatch 86.22 82.91(-3.31) 76.48(-9.74) 78.09(-8.13) 90.35 89.14(-1.21) 91.08(0.73) 92.75(2.40)
DOEPatch(YF) 42.25 66.51(21.26) 90.26(48.01) 88.92(46.67) 47.17 60.61(13.44) 81.10(33.93) 78.85(31.68)
DOEPatch(YYF) 47.86 76.82(28.96) 88.07(40.21) 89.01(41.15) 48.49 66.55(18.06) 82.58(34.09) 82.02(33.53)
IV-G Ablation Studies